OpenStudy (anonymous):

a certain radioactive substance loses one third of its original mass in 5 days. find its half-life using the half-life function N(x) = c(0.5)^(x/h)

OpenStudy (anonymous):

m-m/3=m/2^t t=log_2(1.5) HT=5/t

OpenStudy (anonymous):

why is it m-m/3?

OpenStudy (anonymous):

lost one third

OpenStudy (anonymous):

okay but if i had to use the equation given, would it be N(x)=(1/3)(0.5)^(x/5)

OpenStudy (anonymous):

I don't know what is that !

thats the equation i was given to use ):